A researcher is conducting tests of manual dexterity among college students. The time to complete a certain task is normally distributed with a mean of 30.3 seconds and a standard deviation of 7.2 seconds. Find the probability that a randomly selected student takes more than 28.6 seconds to complete the task
Answer #1

Solution :

Given that ,

P(x > 28.6) = 1 - P(x < 28.6)

= 1 - P[(x - ) / < (28.6 - 30.3) / 7.2)

= 1 - P(z < -0.2361)

= 1 - 0.4067

= 0.5933

Probability = 0.5933